Update ....

John Shott shott at stanford.edu
Sat Sep 6 19:00:01 PDT 2008


Plasma Quest community:

I know that a number of you are anxiously awaiting the return to service 
of the Plasma Quest following its RF and DC bias issues.  I had meant to 
send out an update last night ... and then  got side tracked with air 
compressor stuff.

In any event, we made quite a bit of progress today.  The machine is 
fundamentally re-assembled and we have checked that the chuck can now 
withstand at least 300 volts in atmosphere with no current draw.  We 
believe that this helps to confirm the hypothesis that RF problems were 
created by wafer chips in the space between the chuck and the dark space 
shield.

If things go well tomorrow morning, we will expect to do the final 
alignment of the clamp assembly, hook up the RF supplies, and take 
another look at the gate valve to make sure that it is behaving 
appropriately.

If things proceed smoothly we will do our best to get the system under 
vacuum and to test it operationally.

We'll try to send out another update  tomorrow afternoon.

Have  a good weekend,

John




More information about the pquest mailing list